a cricket ball of mass 150g is moving with a velocity of 12m/s and is hit by a bat so that it turns back with a velocity of 20m/s. The force of blow acts for a time of 0.01s. then find the change in momentum of ball and also find the average force exerted by the bat on the ball.

Grade:11

3 Answers

Arun
6 years ago

Mass of the ball, m = 150 g = 0.15 kg

Initial speed of the ball, u = -12 m/s

Final speed of the bal after being hit, v = 20 m/s

Impulse = change in momentum = final momentum initial momentum

= Impulse = 0.15 × 20 0.15 × (-12) = 4.8 Ns

shifa
5 years ago
We know that Force = Impulse/ Time 
Impulse = mass * change in velocity
Time = 0.01 s
Mass m = 150 g = 1.5 x 10 ̄ ¹ kg
Change in velocity = Final - intial velocity = 20 - (-12) = 32 m/s
Therefore Force = mass * change in velocity / time = 1.5 x 10 ̄ ¹ x 32 / 0.01

= 15 x 32 = 480 N
Average force on the ball is 480 N 
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
Change in momentum of the ball = mv – mu = 0.15 x (20 – ( – 12)) = 0.15 x 32 = 4.8 Ns
From impulse momentum theorem, we know,
Impulse = change in Linear momentum
hence, F.Δt = 4.8
hence, F = 4.8/0.01 = 480 N
 
Hope it helps.
Thanks and regards,
Kushagra